OpenStudy (anonymous):

Write the equation of the line that passes through (-4, 12) and (7, 23). A. y = x + 16 B. y = 3x + 18 C. y = 1/6 x + 16 D. y = 1/3 x + 18

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Equate the slopes defined by each point to each other and solve the result for y.\[\frac{y-23}{x-7}=\frac{y-12}{x+4}\]\[y = x + 16 \]